OpenStudy (annah.bee):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel? ______ mph

OpenStudy (hellokitty17):

Let speed of car B be x Speed of car A is (x - 15) Distance traveled by car B in 1.5 h = x * 1.5 Distance traveled by car A in 2 h = (x - 15) * 2 These two distances are equal Therefore, 1.5x = 2(x - 15) 1.5x = 2x - 30 subtract 1,5x 0 = 0.5x - 30 add 30 30 = 0.5x divide by 0.5 30/0.5 = x x = 60

OpenStudy (hellokitty17):

@anna.baee there you go!

OpenStudy (hellokitty17):

medal plz

OpenStudy (danjs):

A distance is = rate traveling x time Both cars move same distance, rate of A* time of A = rate of B * time of B

OpenStudy (annah.bee):

it's annah bee btw hahaha but thank you